A pharmaceutical company receives large shipments of aspirin tablets. The acceptance sampling plan is to randomly select and test 16 tablets, then accept the whole batch if there is only one or none that doesnt meet the required specifications. If a particular shipment of thousands of aspirin tablets actually has a 5% rate of defects, what is the probability that this whole shipment will be accepted? The probability that this whole shipment will be accepted is Round to three decimal places as needed.)
Answer #1

X ~ Binomial (n,p)

Where n = 16, p = 0.05

Binomial probability distribution is

P(X) = nCx px (1-p)n-x

P( X <= 1) = P( X = 0) + P( X = 1)

= 16C0 0.050 0.9516 + 16C1 0.051 0.9515

= 0.811 (Rounded to three decimal places)